How many C atoms are present in the sample of C3H8 with mass 3.81 g? Avogadro’s number is 6.022 × 1023. Enter your answer using scientific notation and to three significant digits.
Answer #1

Number of moles of C3H8 = 3.81 g/ 44.1 g/mol = 0.0864 mole

1 mole of C3H8 = 6.022*10^23 molecules so

0.0864 mole of C3H8 =

= 0.0864 mole of C3H8 *(6.022*10^23 molecules / 1 mole of C3H8)

= 0.520*10^23 molecules

1 molecule of C3H8 contains 3 Carbon atoms so

0.520*10^23 molecules will contain

= 0.520*10^23 molecules *(3 Carbon atoms / 1 molecule of C3H8)

= 1.56*10^23 atoms

Therefore, the number of C atoms present are 1.56*10^23